Determine the volume in gallons of a 32.5lb mass that has a density of 2.875 g/mL.
1
Expert's answer
2020-11-12T07:10:46-0500

32.5lb to gram = 32.5 × 453.6 = 14742g

Volume = mass / density

Vol= 14742g/ 2.875g/mL

Volume = 5128 mL

5128mL to gallon = 5128/3785 = 1.35 gallon